# Coherence-Mediated Quantum Thermometry in a Hybrid Circuit-QED Architecture

### v1

@@ Appendix D Derivation of Qubit-Only-Based Coherence Envelope and QFI
$C_{q}(\tau)=\exp\!\Big[-i\,2\chi_{a}\bar{n}\,\tau\Big]\;\exp\!\Big[-(2\chi_{a})^{2}\,\bar{n}(\bar{n}+1)\,f_{\kappa}(\tau)\Big],$ (33)
with
$f_{\kappa}(\tau)=\frac{\kappa_{a}\tau-1+e^{-\kappa_{a}\tau}}{\kappa_{a}^{2}}.$ (34)
This expression interpolates smoothly between $f_{\kappa}(\tau)\!\simeq\!\tfrac{1}{2}\tau^{2}$ for $\kappa_{a}\tau\!\ll\!1$ (quasi-static diffusion) and $f_{\kappa}(\tau)\!\simeq\!\tau/\kappa_{a}$ for $\kappa_{a}\tau\!\gg\!1$ (Markovian exponential decay) [35, 33, 1].
From Eq. (33), the amplitude is $C_{q}(\tau,T)=e^{-\Gamma_{q}(T)}$, with
$\Gamma_{q}(T)=(2\chi_{a})^{2}\bar{n}(\bar{n}+1)f_{\kappa}(\tau),$
and the phase is $\Phi_{q}(\tau,T)=2\chi_{a}\tau\bar{n}$. The derivatives are
$\partial_{T}C_{q}=-\,e^{-\Gamma_{q}}\,\Gamma_{q}^{\prime}(T),\qquad\Gamma_{q}^{\prime}(T)=(2\chi_{a})^{2}f_{\kappa}(\tau)\,\big(1+2\bar{n}(T)\big)\,\partial_{T}\bar{n}(T),$
and
$\partial_{T}\Phi_{q}(\tau,T)=2\chi_{a}\,\tau\,\partial_{T}\bar{n}(T).$ (35)
Substituting into Eq. (13), the QFI becomes
$\mathcal{F}_{q}=\frac{(\partial_{T}\Gamma_{q})^{2}}{e^{2\Gamma_{q}}-1}+e^{-2\Gamma_{q}}\big(2\chi_{a}\tau\big)^{2}\,\big(\partial_{T}\bar{n}\big)^{2}.$ (36)
[…]
Importantly, in this protocol the qubit itself is the sensor: it accumulates the thermal information through the same dispersive coupling that simultaneously dephases it. This trade-off is intrinsic to the qubit-only scheme and cannot be eliminated by echo or refocusing, since the stochastic diffusion encodes the very information being extracted. Thus, practical operation requires choosing $\tau\approx\tau^{\star}$ and balancing the gain from large $\chi_{a}$ against the inevitable loss of visibility. In contrast, the coherence-mediated scheme off-loads the sensing burden onto a high-$Q$ probe mode: while the probe collects the temperature-dependent phase diffusion, the qubit can be dynamically decoupled from it, enabling much longer interaction times and leveraging the stability of a 3D cavity to preserve visibility.
Appendix E Sixth–order perturbation theory for the two–qubit mediated cross–Kerr
[…]
$H_{0}/\hbar=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+\tfrac{\omega_{1}}{2}\sigma_{z1}+\tfrac{\omega_{2}}{2}\sigma_{z2},$
and the interaction term
$V/\hbar=g_{a1}(a\sigma_{1}^{+}+a^{\dagger}\sigma_{1}^{-})+g_{b2}(b\sigma_{2}^{+}+b^{\dagger}\sigma_{2}^{-})+J_{XY}(\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}).$
@@ Appendix E Sixth–order perturbation theory for the two–qubit mediated cross–Kerr
$H_{\mathrm{eff}}\supset\lambda\,a^{\dagger}a\,b^{\dagger}b,\qquad\lambda\;\sim\;\frac{g_{a1}^{2}g_{b2}^{2}J_{XY}^{2}}{\Delta_{a1}^{2}(\Delta_{a1}+\Delta_{12})^{2}(\Delta_{a1}+\Delta_{b2})}\;+\;\text{(mirror and permutations)}.$
Additionally, there are other time orderings permitted by Rayleigh–Schrödinger perturbation theory. Summing all allowed ladders restores a symmetric expression in $(\Delta_{a1},\Delta_{b2},\Delta_{12})$, and gives the effective cross Kerr
$\lambda=\frac{8\chi_{a1}\chi_{b2}J_{XY}^{2}}{\Delta_{12}^{3}},$ (38)
with $\chi_{a1}=-g_{a1}^{2}/\Delta_{a1}$ and $\chi_{b2}=-g_{b2}^{2}/\Delta_{b2}$.
@@ Appendix F Parasitic Dispersive Couplings and Qubit Dephasing
While the cross-Kerr interaction between the probe ($b$) and thermal ($a$) modes is the desired sensing channel, the qubit also has residual dispersive couplings to each mode, with strengths $\chi_{b}$ and $\chi_{a}$. These couplings do not encode thermometric information; they act as parasitic dephasing pathways that reduce Ramsey visibility during the readout window of duration $\tau_{R}$.
If the qubit couples dispersively to a bosonic mode $m$ with photon-number variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$, a second-order cumulant expansion gives the coherence envelope during the Ramsey window [35]
$C_{q}^{(m)}(\tau_{R})=\exp\!\left\{-\,4\chi_{m}^{2}\,\mathrm{Var}(n_{m})\left[\frac{\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}}}{\kappa_{m}^{2}}\right]\right\}.$ (39)
Equation (39) smoothly interpolates between Gaussian decay in the quasi-static limit ($\kappa_{m}\tau_{R}\ll 1$) and exponential decay in the Markovian limit ($\kappa_{m}\tau_{R}\gg 1$).
For the high-$Q$ 3D cavity probe, $\kappa_{b}\tau_{R}\ll 1$, so (39) reduces to
$C_{q}^{(b)}(\tau_{R})\approx\exp\!\left[-2\chi_{b}^{2}\,\mathrm{Var}(n_{b})\,\tau_{R}^{2}\right],$ (40)
with photon-number variance for a displaced thermal state
$\mathrm{Var}(n_{b})=\bar{n}_{b}(1+\bar{n}_{b})+|\alpha|^{2}(1+2\bar{n}_{b}).$
Here $|\alpha|^{2}$ is the coherent population and $\bar{n}_{b}$ the residual thermal occupation. Increasing $|\alpha|$ boosts temperature-to-phase transduction in the probe but also increases parasitic dephasing during the Ramsey readout via $\chi_{b}$. The information encoded in the probe during the imprinting interval is independent of $\chi_{b}$; $\chi_{b}$ primarily sets the speed and backaction of the readout.
Qubit dephasing due to residual coupling to the thermal mode also follows Eq. (39), with
$\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a}).$
In practice the thermal mode is implemented as a low-$Q$ resonator with $\kappa_{a}\tau_{R}\gtrsim 1$, placing Eq. (39) in the Markovian regime. Expanding for $\kappa_{a}\tau_{R}\gg 1$ yields
$C_{q}^{(a)}(\tau_{R})\approx\exp\!\left[-\,\frac{4\chi_{a}^{2}\,\mathrm{Var}(n_{a})}{\kappa_{a}}\,\tau_{R}\right],$ (41)
i.e., an exponential decay with rate $\gamma_{\varphi}^{(a)}=4\chi_{a}^{2}\,\mathrm{Var}(n_{a})/\kappa_{a}$. Thus, a strongly damped thermal resonator acts as a broadband Markovian bath: while $\mathrm{Var}(n_{a})$ grows with temperature, the large linewidth $\kappa_{a}$ suppresses the overall dephasing rate. This contrasts with the high-$Q$ limit, where quasi-static fluctuations induce Gaussian-in-time visibility loss.
In summary, both $\chi_{b}$ and $\chi_{a}$ enter through the general law (39), but with distinct photon statistics and dynamical regimes. Engineering small residual couplings ensures that visibility is limited only during the short Ramsey window, while the probe amplitude $\alpha$ can be used as a controlled resource to optimize sensitivity without incurring unnecessary readout dephasing.
@@ Appendix G Phase-to-Quadrature Mapping
In the coherence-mediated scheme, the coherent probe accumulates a phase shift $\phi_{b}=\lambda\!\int_{0}^{\tau}n_{a}\,dt$ during its interaction with the thermal mode over the interaction time $\tau$. However, its photon number remains fixed, $n_{b}=|\alpha e^{i\phi_{b}}|^{2}=|\alpha|^{2}$, so a plain Ramsey evolution under the dispersive coupling $\chi_{b}\hat{n}_{b}\hat{\sigma}_{z}$ is phase blind.

### v2

@@ Appendix D Derivation of Qubit-Only-Based Coherence Envelope and QFI
$C_{q}(\tau)=\exp[-i\chi_{a}\bar{n}_{a}\,\tau]\,\exp\!\left[-\chi_{a}^{2}\,\bar{n}_{a}(\bar{n}_{a}+1)\,f_{\kappa}(\tau)\right],$ (31)
with $f_{\kappa}(\tau)=(\kappa_{a}\tau-1+e^{-\kappa_{a}\tau})/\kappa_{a}^{2}.$ This expression interpolates between the quasi-static regime $\kappa_{a}\tau\ll 1$, where $f_{\kappa}(\tau)\approx\tfrac{1}{2}\tau^{2}$, and the Markovian regime $\kappa_{a}\tau\gg 1$, where $f_{\kappa}(\tau)\approx\tau/\kappa_{a}$.
From Eq. (31), the coherence amplitude is $|C_{q}(\tau,T)|=e^{-\Gamma_{q}}$, with $\Gamma_{q}=\chi_{a}^{2}\,\bar{n}_{a}(\bar{n}_{a}+1)\,f_{\kappa}(\tau),$ and the accumulated phase is $\Phi_{q}(\tau,T)=\chi_{a}\bar{n}_{a}\,\tau$. Differentiation yields $\partial_{T}\Gamma_{q}=\chi_{a}^{2}f_{\kappa}(\tau)\,(1+2\bar{n}_{a})\,\partial_{T}\bar{n}_{a}$ and $\partial_{T}\Phi_{q}=\chi_{a}\tau\,\partial_{T}\bar{n}.$
Substituting these into the pure-state QFI expression gives
$\mathcal{F}_{q}=\frac{(\partial_{T}\Gamma_{q})^{2}}{e^{2\Gamma_{q}}-1}\;+\;e^{-2\Gamma_{q}}\,\big(\chi_{a}\tau\big)^{2}\,(\partial_{T}\bar{n}_{a})^{2}.$ (32)
Figure 7 illustrates the resulting behavior. For a representative dispersive coupling $\chi_{a}/2\pi=50~\mathrm{kHz}$, the qubit-only QFI $\mathcal{F}_{q}$ exhibits the same qualitative trends as the CM scheme in the quasi-static limit ($\kappa_{a}\tau\ll 1$): the optimal $\mathcal{F}_{q}$ and corresponding interaction time $\tau^{\star}$ both decrease with increasing bath temperature $T$, as shown in Fig. 7(a). Fig. 7(b) shows that in the quasi-static regime (small $\kappa_{a}$) the competition between information gain and dephasing yields a finite optimal interaction time $\tau^{\star}$, whereas for large $\kappa_{a}$ the dynamics become effectively Markovian and $\mathcal{F}_{q}(\tau)$ increases monotonically with $\tau$. As shwon in Fig. 7(c), stronger dispersive couplings $\chi_{a}$ drive the optimum to shorter times $\tau^{\star}$, but the maximum achievable $\mathcal{F}_{q}$ remains nearly independent of $\chi_{a}$, indicating that stronger coupling primarily compresses the optimal sensing window rather than enhancing the overall per-shot information.
Importantly, in the QB scheme the qubit is both sensor and memory: the same dispersive coupling that encodes temperature also dephases the qubit. This intrinsic trade-off cannot be removed by echo without erasing the signal, in contrast to the coherence-mediated scheme where the temperature-dependent diffusion is absorbed entirely by the probe mode. Thus, practical operation requires choosing $\tau\approx\tau^{\star}$ and balancing the gain from large $\chi_{a}$ against the inevitable loss of visibility. In contrast, the CM scheme off-loads the sensing burden onto a high-$Q$ probe mode: while the probe collects the temperature-dependent phase diffusion, the qubit can be dynamically decoupled from it, enabling much longer interaction times and leveraging the stability of a 3D cavity to preserve visibility.
Appendix E Sixth-order perturbation theory for the two–qubit mediated cross–Kerr
To make the virtual process transparent, we evaluate the cross-Kerr coupling between bosonic modes $a$ and $b$ using straightforward sixth–order perturbation theory (PT) [30]. Two transmon qubits $Q_{1}$ and $Q_{2}$ bridge the modes in series ($a$-$Q_{1}$-$Q_{2}$-$b$).
The interaction term of the system Hamiltonian is
$V/\hbar=\,$ $g_{a1}(a\sigma_{1}^{+}+a^{\dagger}\sigma_{1}^{-})+g_{b2}(b\sigma_{2}^{+}+b^{\dagger}\sigma_{2}^{-})\;+$
$J_{XY}(\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}).$
@@ Appendix E Sixth-order perturbation theory for the two–qubit mediated cross–Kerr
$\lambda\;\sim\;\frac{g_{a1}^{2}g_{b2}^{2}J_{XY}^{2}}{\Delta_{a1}^{2}(\Delta_{a1}+\Delta_{12})^{2}(\Delta_{a1}+\Delta_{b2})}.$
Additionally, there are other time orderings permitted by Rayleigh–Schrödinger perturbation theory. Summing all allowed ladders restores a symmetric expression in $(\Delta_{a1},\Delta_{b2},\Delta_{12})$, and gives the effective cross-Kerr
$\lambda=\frac{8\chi_{a1}\chi_{b2}J_{XY}^{2}}{\Delta_{12}^{3}}\,\Big[1+\mathcal{O}\!\big(\tfrac{\Delta_{12}}{\Delta_{a1}},\;\tfrac{\Delta_{12}}{\Delta_{b2}}\big)\Big],$ (33)
with $\chi_{a1}=-g_{a1}^{2}/\Delta_{a1}$ and $\chi_{b2}=-g_{b2}^{2}/\Delta_{b2}$. Here, the subleading terms account for finite anharmonicity and detuning sign [4, 21].
In practice, the same cross–Kerr interaction can be gated by flux-tuning one of the bridge qubits (or the intermediate coupler) to temporarily reduce the effective detuning $\Delta_{12}$ during the sensing window. Such a flux-tunable transmon implementation enables transiently enhanced $J_{XY}/\Delta_{12}$ while the system remains dispersive in its idle configuration, thereby enhancing the cross–Kerr rates without compromising qubit coherence.
It is noteworthy that the same cross-Kerr interaction that enables temperature sensing also mediates measurement backaction: fluctuations of the probe photon number $\hat{n}_{b}$ induce an additional dephasing rate in the absorber, $\Gamma_{\phi,a}\propto\lambda^{2}\bar{n}_{b}/\kappa_{b}$, which manifests as extra decay in the probe’s coherence envelope during the interaction [50]. Minimizing this backaction constrains the allowable probe photon number $\bar{n}_{b}=\alpha^{2}$ and defines the optimal balance between signal transduction and absorber disturbance. In a gated implementation, a flux–tunable bridge can be activated only during the sensing window, thereby limiting the accumulated backaction while preserving the desired transient cross-Kerr strength.
@@ Appendix F Parasitic Dispersive Couplings and Qubit Dephasing
While the cross-Kerr interaction between the probe ($b$) and thermal ($a$) modes is the desired sensing channel, the qubit also retains residual dispersive couplings to each mode, with strengths $\chi_{b}$ and $\chi_{a}$. These couplings do not encode thermometric information; instead, they generate parasitic dephasing during the Ramsey readout window of duration $\tau_{R}$.
If the qubit couples dispersively to a bosonic mode $m$ with photon-number variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$, a second-order cumulant expansion gives the general coherence envelope during the Ramsey window [25]:
$C_{q}^{R}(\tau_{R})=\exp\!\left[-\,4\chi_{m}^{\,2}\,\mathrm{Var}(n_{m})\,f^{R}(\kappa_{m},\tau_{R})\right],$ (34)
with $f^{R}(\kappa_{m},\tau_{R})=(\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}})/\kappa_{m}^{2}.$ This expression captures the full crossover between slowly fluctuating and rapidly fluctuating photon-number noise in a single closed form.
For the thermal mode, linewidth $\kappa_{a}$ and temperature-dependent variance $\mathrm{Var}(n_{a})$ jointly determine the associated dephasing rate, but this pathway plays no constructive role in sensing and should be mitigated by engineering $\chi_{a}$ to be small.
For the probe mode, which is initialized in a displaced thermal state, the photon-number variance is $\mathrm{Var}(n_{b})=\bar{n}_{b}(1+\bar{n}_{b})+|\alpha|^{2}(1+2\bar{n}_{b}),$ Increasing $|\alpha|$ enhances temperature-to-phase transduction in the probe but also increases parasitic qubit dephasing through the residual coupling $\chi_{b}$ during the short Ramsey readout.
Therefore, applying a Hahn echo during the readout window yields a modified coherence envelope that suppresses quasi-static photon-number fluctuations in the probe. For dispersive coupling to mode $b$, the exact echo envelope is [24, 33, 14]
$C_{q}^{E}(\tau_{R})=\exp\!\left[-\,4\chi_{b}^{\,2}\,\mathrm{Var}(n_{b})\,f^{E}(\kappa_{b},\tau_{R})\right],$ (35)
with $f^{E}(\kappa_{b},\tau_{R})=(\kappa_{b}\tau_{R}-3+4e^{\kappa_{b}\tau_{R}/2}-e^{-\kappa_{b}\tau_{R}})/\kappa_{b}^{2}.$
All parasitic pathways enter through the unified coherence laws in Eqs. (34)-(35), with distinct photon statistics for the probe and thermal modes. Engineering small $\chi_{b}$ and $\chi_{a}$ ensures that the qubit maintains high visibility during the brief Ramsey readout, while the probe amplitude $|\alpha|$ and the cross-Kerr interaction remain free parameters for optimizing thermometric sensitivity. Echo sequences can be used to suppress quasi-static probe-induced fluctuations arising from the high-$Q$ probe cavity.
@@ Appendix G Phase-to-Quadrature Mapping
In the CM scheme, the coherent probe accumulates a phase shift $\phi_{b}=\lambda\!\int_{0}^{\tau}n_{a}\,dt$ during its interaction with the thermal mode over the interaction time $\tau$. However, its photon number remains fixed, $n_{b}=|\alpha e^{i\phi_{b}}|^{2}=|\alpha|^{2}$, so a plain Ramsey evolution under the dispersive coupling $\chi_{b}\hat{n}_{b}\hat{\sigma}_{z}$ is phase blind.